OpenStudy (anonymous):

\[{1+i}\over{-4-8i}\]

OpenStudy (anonymous):

multiply and divide with -4+8i

OpenStudy (anonymous):

what??

hartnn (hartnn):

to make the denominator real,u need to multiply and divide by the complex conjugate of the denominator....u know whats complex conjugate of -4-8i ??

OpenStudy (anonymous):

oh okay i remember now

hartnn (hartnn):

then u can use the formula for denominator: (a+bi)(a-bi)=a^2+b^2

OpenStudy (anonymous):

what is i^2, again?

hartnn (hartnn):

i^2 = -1

OpenStudy (anonymous):

no. (-4-8i)(-4+8i)=(-4)^2-(8i)^2=16-64i^2=16-64(-1)=16+64=80

OpenStudy (anonymous):

\[{{12i+3}\over{80}}~~~~~?\]

OpenStudy (anonymous):

the question says to write it in standard form....???

hartnn (hartnn):

(i-3)/20

OpenStudy (anonymous):

in denominator 80 should be there in numerator you should have (i+1)(-4+8i) after multiplication you should have. -4i+8i^2-4+8i=-12+4i this is the numerator.

OpenStudy (anonymous):

no its (1+i)(4+8i)

OpenStudy (anonymous):

because it was (-4-8i) so the complex conj. is (4+8i)

hartnn (hartnn):

nopes,for complex conjugate u flip the sign of imaginary part only!

OpenStudy (anonymous):

the conjugate does not change the real part so sonjugate should be -4+8i

OpenStudy (anonymous):

oh okay right...so its\[{-12i+3}\over80\]

OpenStudy (anonymous):

but how do i write that in standard form??

OpenStudy (anonymous):

\[4i-12\over80\] is what im getting now

OpenStudy (anonymous):

which is\[i-3\over20\]

OpenStudy (anonymous):

in the numerator it should be -12+4i standard form is a+bi so (-12+4i)/(80) =-12/80+4i/80 =-3/20+1/20 i here a=-3/20 and b=1/20 =-3/20+1/20 i is standard form.

OpenStudy (anonymous):

(i-3)/20 is standard form?

OpenStudy (anonymous):

you can say that. but usually complex no are expressed in z=a+ib if you expand it you will have =-3/20+i/20

hartnn (hartnn):

do u have choices,if not -3/20+i/20 is standard form....

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

welcome :)